Given the function G(w)=|w|-1 find each of the following:
1. G(-54)=
2. G(53)=
3. G(3/4)=
4. G(-3)=

Answer by drk(1908)

1. G(-54)= abs(-54) - 1 = 54 - 1 = 53
2. G(53)= abs(53) - 1 = 53 - 1 = 52
3. G(3/4)=abs(3/4) - 1 = 3/4 - 1 = -1/4
4. G(-3)=abs(-3) - 1 = 3 - 1 = 2